OpenStudy (anonymous):

Expand the following equations. -2(3x-5) 1/5(3x+1/2) 0.5(5x-2.2)

OpenStudy (bibby):

distribute the factor to each term on the inside recall that \(a(b+c)=a*b+a*c\) and \(a(b-c)=a*b-a*c\)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

GJAVCUASZVFUESFVUFVIAYFJED

OpenStudy (anonymous):

Im sorry that was my little bad brother